Differentiate the following:
5ln(2x−2y)−2y^2+5x=0
1
Expert's answer
2019-06-21T11:42:37-0400
"F=5ln(2x-2y)-2y^2+5x"

In calculus, a method called implicit differentiation makes use of the chain rule to differentiate implicitly defined functions

"y'=-\frac{F'_x}{F'_y}"

take the derivative with respect to x from F

"{F'_x}=5ln(2x-2y)dx-2y^2dx+5xdx"



"(5ln(2x-2y))'dx=5(ln(2x-2y)'dx"


"(u(v))'=u'(v)v'"

"5(ln(2x-2y)'dx=5\frac{(2x-2y)'dx}{2x-2y}=5\frac{2}{2x-2y}"


"(2y^2)'dx=0"




"(5x)'dx=5"




"{F'_x}==5\frac{2}{2x-2y}+5"

take the derivative with respect to y from F

"{F'_y}=5ln(2x-2y)dy-2y^2dy+5xdy"

"(5ln(2x-2y)'dy=5(ln(2x-2y)'dy=5\frac{(2x-2y)'dy}{2x-2y}=5\frac{-2}{2x-2y}"


"(2y^2)'dy=4y"


"(5x)'dy=0"





"{F'_y}==5\frac{-2}{2x-2y}-4y"


"{F'_x}=\frac{5}{x-y}+5=5\frac{1+x-y}{x-y}"


"{F'_y}=-(\frac{5}{x-y}+4y)=-(\frac{5+4y(x-y)}{x-y})"


"y'=-\frac{5\frac{1+x-y}{x-y}}{-(\frac{5+4y(x-y)}{x-y})}"


"y'=\frac{5\frac{1+x-y}{x-y}}{\frac{5+4y(x-y)}{x-y}}""y'=\frac{5(1+x-y)}{5+4y(x-y)}"